# Compensation behaviour in trilayered anisotropic 6-state clock model

### v2

@@ Abstract
The equilibrium behaviours of a trilayered 6-state clock model has been investigated by Monte Carlo simulation. The intralayer interaction is considered ferromagnetic, whereas the interlayer interaction is antiferromagnetic. Periodic boundary conditions(PBC) are applied in XY plane and open boundary condition (OBC) is applied along Z-direction. The thermodynamic behaviours of sublattice magnetisation, total magnetisation, magnetic susceptibility and the specific heat are studied as functions of the temperature. The interesting compensation phenomenon, i.e., the vanishing of the total magnetisation even with nonvanishing sublattice magnetisation, has been observed. The compensation temperature and the critical temperature are found to depend on the strength of single-site anisotropy. The comprehensive phase diagram is shown in the temperature-anisotropy plane. The finite size analysis has been done through the temperature dependence of fourth order Binder cumulant, and the scaling exponents are estimated.
@@ 1 Introduction
The ferromagnetic models showing the unusual ordered phase even in absence of conventional long-range ferromagnetic order is a remarkable discovery[1, 2]. The continuous symmetric (SO(2)) planar ferromagnetic model shows this behaviour in two dimensions with vortex-antivortex special kind of ordered state[3, 4, 5]. The unconventional metastable behaviour and phase transition in this kind of model has drawn the attention of the researchers in last four decades. The celebrated BKT transition is now under the spotlight of modern research in condensed matter physics and statistical mechanics.
Is this kind of unusual phase the property of continuous symmetry only? This question has triggered to investigate the discrete symetric q-state clock model, where the continuous symetric XY model can be restored in the limit $q\to\infty$. The properties of q-state clock model has been investigated[6] for q=4, 5 and 6 in two dimensions by Monte Carlo renormalization group analysis to study the phase transitions. The Kosterlitz-Thouless type intermediate phase has been found[7] in two dimensional 6-state clock model by Monte Carlo simulations. The KT type phase transition has been found to occur in the range between ${{k_{B}T}\over{J}}=0.68\pm 0.02$ and ${{k_{B}T}\over{J}}=0.92\pm 0.01$. However, it is confirmed later that there exists a Kosterlitz-Thouless (KT) like phase in the ferromagnetic 6-clock model on the square lattice through studies of the interfacial free energy estimated by Monte Carlo simulations and the authors estimated[8] that lower and upper transition temperatures are 0.75 and 0.90 respectively. Interestingly, in a recent study[9], the machine-learning detection of BKT transition was found in q-state clock model with the neural network based estimates of the two critical points as ${{k_{B}T}\over{J}}=0.410$ and ${{k_{B}T}\over{J}}=0.921$.
The classical 2D clock model is known to have a critical phase with Berezinskii–Kosterlitz–Thouless (BKT) transitions. These transitions have logarithmic corrections which make numerical analysis difficult. Hence, the classical 2D six-state clock model has been studied [10] by mapping it to the quantum self-dual 1D six-state clock model and confirmed that the self-dual point has a precise numerical agreement with the analytical result. It has also been argued that the degeneracy of the excited states at the self-dual point from the effective field theoretical point of view.
@@
The thermodynamic behaviours of clock model in layered structured lattices are not yet reported in the well developed literature of this specific model. However, the behaviours of layered Ising and Blume-Capel models are well studied. Let us briefly mention a few selected studies on layered Ising and Blume-Capel models. The main interesting phenomenon in the thermodynamic behaviours of layered Ising and Blume-Capel models is the compensation transition. The compensation temperatures and magnetic susceptibility of a mixed ferro-ferrimagnetic ternary alloy have been studied[16] by Monte Carlo simulations and observed outstanding behaviour of magnetic susceptibility near the compensation and critical temperature. The compensation temperature has been found[17] to depend of the anisotropy in a mixed spin (spin-${1}\over{2}$} and spin-$1$) Heisenberg ferrimagnet on a simple cubic lattice, studied by Oguchi approximation. The anisotopy dependent spin compensation temperature has also been noticed[18] in mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) ferrimagnetic Ising model studied by mean field approximation. In this case, the possibility of multiple compensation temperature was indicated. The impurity- concentration dependent compensation[19] and the anisotropy dependent compensation[20] was also noticed in the trilayered Ising and Blume-Capel models, respectively. The meanfield and effective field theory has been employed [21] to study the interaction dependent compensation in spin-${1}\over{2}$} Ising trilayered ferrimagnetic system. The phase diagram (compensation and no-compensation) was drawn in parameter space. The antiferromagnetic systems showed the compensation behaviours in Ising mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) core/shell nanowire[22] and in mixed spin (spin-${7}\over{2}$} and spin-1{1}) antiferromagnetic ovalene nanowire[23]. The compensation wad found[24] Ising mixed spin (spin-${3}\over{2}$} and spin-${1}\over{2}$}) graphene bilayered ferrimagnet. The effects of exchange interactions, crystal field and external magnetic field on the compensation temperature were studied [24] by Monte Carlo simulation. The evidence of engineering the compensation in ferrimagnetic iron garnet by tuning the ratio of in-plane and out of plane magnetisations can be found in [25]. An interesting spin compensation and first-order reentrant behaviour has been observed[26] in mixed spin (spin-1 and spin-2) Ising ferrimagnetic system studied by Monte Carlo simulation.
@@ 2 Model and numerical details
We consider a classical trilayered 6-state clock model defined on a simple cubic lattice, where each of the three monoatomic layers consists exclusively of either type-A or type-B atoms. The magnetic interactions and anisotropic contributions to the energy of the system are captured by the following Hamiltonian:
$\mathcal{H}=-J_{ferro}^{intra}\sum_{\langle ij\rangle}\vec{S_{i}}\cdot\vec{S_{j}}-J_{antiferro}^{inter}\sum_{\langle ij\rangle}\vec{S_{i}}\cdot\vec{S_{j}}-D\sum_{i}[({S_{i}}^{x})^{2}-({S_{i}}^{y})^{2}]$ (1)
@@
The first term in Equation-1 accounts for the intralayer ferromagnetic interactions between nearest neighbour spins $\langle ij\rangle$ within the same atomic plane, with a positive coupling exchange $J_{ferro}^{intra}>0$. The second term represents interlayer antiferromagnetic coupling between the adjacent layers, involving nearest neighbour sites $\langle ij\rangle$ across the layers, governed by a coupling constant $J_{antiferro}^{inter}<0$. The third term introduces an anisotropy D, which destroys the discrete rotational symmetry in the $xy$-plane and favours the spin-alignment in specific direction. Positive values of $D$ prefer specific spin orientations, thereby reinforcing the discrete six-fold symmetry and enhancing the stability of the selected clock states. We impose periodic boundary conditions (PBC) in the x and y directions within each layer and along the stacking direction z; we adopt free boundary conditions to capture the finite thickness of the trilayered structure.
To study the compensation behaviour in the trilayered 6-state clock model, we employ the Monte Carlo simulation technique based on the Metropolis algorithm with random single-spin updates. The system consists of three stacked square lattices, each consisting of $L\times L$ sites, resulting in a total of $3L^{2}$ spins. For the majority of our simulations, we consider a system size of $L=128$, corresponding to $128\times 128$ spins per layer. In order to examine finite-size effects, additional simulations are performed for various lattice sizes. Initially, spin configurations are generated at high temperature with randomly assigned orientations chosen from the six allowed states in the clock model. Then the system is allowed to evolve using Metropolis updating scheme: during each Monte Carlo steps (MCS), $L^{2}$ spin-flip attempts are made, where a trial spin selected at random and updated according to the Metropolis acceptance probability[27]: $P_{f}=Min[\exp{-\frac{\delta E}{k_{B}T}},1]$. $\Delta E$ is the change in energy associated with the proposed spin flip, and $k_{B}(=1)$ is the Boltzmann constant. The simulation spans a duration $3\times 10^{5}$ MCSS within an initial transient segment of $1.5\times 10^{5}$ MCSS excluded to achieve the assumed ergodic limit. To characterize the thermodynamic behaviour of the 6-state clock model, we calculate several key observables that capture the collective magnetic ordering of the system. The sublattice magnetisation provides insight into the local symmetry breaking and is defined for each sublattice $k\in{1,2,3}$ as the average spin components in the x and y directions:
@@
where $\langle E\rangle$ is the average energy density calculated from the Hamiltonian. We also calculate the thermal variation of fourth-order Binder cumulant for different system sizes,
$U_{L}=1-\frac{\langle M^{4}\rangle}{3\langle M^{2}\rangle^{2}}$
@@ 3.1 Isotropic regime:$D=0$
To visualise the spin ordering, we also examine the lattice morphology at representative temperatures. Spin configurations snapshots displayed in Figure-3 reveal a transition from a disordered paramagnetic state at high temperature to a well-aligned domain at low temperature for different layers. At very high temperature $T>>T_{c}$ spin are randomly oriented in all three layers. Six colours denote the six specific clock states. For the isotropic regime $(D=0)$, at high temperature, all six directions are equally preferable. As temperature decreases, the spins at each layer progressively align, leading to the nucleation and growth of the magnetic domain, which marks the onset of long-range order and results a non-zero total magnetisation. Near $T_{comp}$, distinct domain patterns emerge with opposing magnetisation contributions, reflecting the layer-wise compensation behaviour. To quantitatively validate the spin morphology, we analysed the angular distribution of spins across the three layers at different temperatures. The normalised spin population (fraction of total spins) is plotted as a function of six discrete clock states ( corresponding spin angle 0, $\pi/3$, $2\pi/3$, $3\pi/3$, $4\pi/3$, $5\pi/3$) for each layer. At high temperature, the spin population is nearly uniform across all angles in all three layers, reflecting an isotropic regime where all spin states are equally probable. As temperature decreases towards the critical value, certain spin angles become more populated, indicating the onset of preferred alignment. Near compensation temperature, layer 1 and 3 exhibit a dominant population at $\theta=\pi/3(n=2)$, whereas layer 2 prefers $\theta=5\pi/3(n=6)$. Importantly, the other spin angles are populated such that net magnetisation of the whole system becomes zero, while each layer individually retains a finite magnetisation. Upon further cooling to low temperature, layers 1 and 3 show nearly complete alignment at $\theta=\pi/3$ while layer 2 remains oriented in $5\pi/3$. The visual and quantitative results together confirm the layered character of magnetic ordering and the competing interactions that drive the compensation phenomenon.
@@ 3.2 Anisotropic regime: $D>0$
To further elucidate the role of anisotropy in shaping the spin morphology, we also examined the angular distribution of spins for the case of $D=2.0$, where the easy-axis energetically favours spin orientations along $\theta=0$ or $\theta=\pi$. This is shown in Figure-9. At high temperature, the majority of spins in all layers are distributed near $\theta=0$, and $\theta=\pi$ reflecting the dominant easy-axis alignment imposed by the positive anisotropy. Upon cooling, the distribution becomes progressively more ordered, with the relative populations $\theta=0$ and $\theta=\pi$ starting to compete. Near the compensation temperature, a distinct redistribution occurs: layers 1 and 3 exhibit a stronger spin population at $\theta=\pi$, whereas layer 2 retains it alignment at $\theta=0$. This antiparallel arrangement of sublattices leads to a vanishing total magnetisation while preserving finite sublattice magnetisation. At very low temperature, the system reaches a nearly complete alignment where layers 1 and 3 are oriented along $\theta=\pi$ and layer 2 aligned along $\theta=0$. At sufficiently low temperatures, the system recovers a finite total magnetisation as thermal fluctuations subside. Thus, while both the isotropic ($D=0$) and anisotropic ($D=2.0$) systems exhibit a compensation behaviour governed by the redistribution of opposing spin populations, the positive anisotropy merely selects preferred axes ($\theta=0,\pi$) without changing the underlying mechanism.
3.3 Finite size analysis
@@
Finite-size effects play a crucial role in accurately characterising the thermodynamic phase transitions in numerical simulations. To investigate the finite-size behaviour of the clock model in anisotropic regime ($D=2.0$), we performed simulations for different system sizes $L=32,64,128,256$. The fourth-order Binder-cumulant $U_{L}(T)$, magnetisation $M(T)$ and corresponding susceptibility $\chi(T)$ are studied as a function of temperature. The numerical results are shown in Figure-10 for the considered system sizes. Figure-10(a) illustrates the Binder Cumulant $U_{L}(T)$ where a well-defined intersection provides a robust indicator of the critical temperature, as it is largely independent of $L$ near the thermodynamic transition. Figure-10(b) shows the magnetisation, revealing the expected order-disorder behaviour as the temperature crosses the transition. The susceptibility curves for different system sizes manifest two salient features: a prominent peak at the thermodynamic critical temperature and a secondary peak at the compensation point. From the Binder cumulant analysis, the intersection point of different $L$ curves yields a critical temperature $T_{c}=1.855$ marking the thermodynamic phase transition at anisotropy $D=2.0$. The susceptibility peak at $T_{c}$ exhibits a systematic increase with increasing system sizes, confirming the growth of critical correlation for the second-order phase transition. The susceptibility diverges in the thermodynamic limit ($L\rightarrow\infty$) according to the scaling $\chi_{max}\sim L^{\frac{\gamma}{\nu}}$. In Figure-11, we present maxima of susceptibility as function of system sizes in logarithmic scale. The straight line fit estimates the critical exponent ratio $\frac{\gamma}{\nu}=0.532$. In contrast, the secondary peak associated with the compensation point does not show such scaling behaviour; its height remains nearly constant across the studied system sizes. This suggests that the compensation phenomena is not a critical transition, but rather a finite, non-divergent reorganisation of the magnetic sublattices. The finite-size analysis confirms that the clock model undergoes a thermodynamic phase transition at $T_{c}$ , where critical fluctuations become system-size dependent, while the compensation point reflects a non-critical feature, inherent to the system’s response.

### v4

# Compensation in trilayered anisotropic 6-state clock model

@@ Abstract
The equilibrium behaviours of a trilayered 6-state clock model have been investigated by Monte Carlo simulation. The intralayer interaction is considered ferromagnetic, whereas the interlayer interaction is antiferromagnetic. Periodic boundary conditions(PBC) are applied in XY plane and open boundary condition (OBC) is applied along Z-direction. The thermodynamic behaviours of sublattice magnetisation, total magnetisation, magnetic susceptibility and the specific heat are studied as functions of the temperature. The interesting compensation phenomenon, when the total magnetisation vanishes keeping the nonzero sublattice magnetisations, has been observed. The compensation temperature and the critical temperature are found to depend on the strength of single-site anisotropy. The comprehensive phase diagram is shown in the temperature-anisotropy plane. The larger systems show the growth of the height of the peak of the susceptibility near the critical temperature.
@@ 1 Introduction
The ferromagnetic models showing the unusual ordered phase even in absence of conventional long-range ferromagnetic order is a remarkable discovery[1, 2]. The continuous symmetric (SO(2)) planar ferromagnetic model shows this behaviour in two dimensions with vortex-antivortex special kind of ordered state[3, 4]. The unconventional metastable behaviour and phase transition in this kind of model has drawn the attention of the researchers in last four decades. The celebrated Berezinskii-Kosterlitz-Thouless (BKT)[5] transition is now under the spotlight of modern research in condensed matter physics and statistical mechanics. It is a transition from bound vortex-antivortex pairs at low temperatures to unpaired vortices and anti-vortices at some critical temperature. The BKT phase gives rise to a special kind of phase where the conventional long range ferromagnetic order is absent. This is observed in two dimensions of SO(2) symmetric models. The planar ferromagnetic XY model in two dimensions shows BKT transitions. The two dimensional superconductor[6], liquid crystalline systems [7] also shows BKT transitions.
Is this kind of unusual phase the property of continuous symmetry only? This question has triggered to investigate the discrete symetric q-state clock model, where the continuous symetric XY model can be restored in the limit $q\to\infty$. The properties of q-state clock model has been investigated[8] for q=4, 5 and 6 in two dimensions by Monte Carlo renormalization group analysis to study the phase transitions. The Kosterlitz-Thouless type intermediate phase has been found[9] in two dimensional 6-state clock model by Monte Carlo simulations. The KT type phase transition has been found to occur in the range between ${{k_{B}T}\over{J}}=0.68\pm 0.02$ and ${{k_{B}T}\over{J}}=0.92\pm 0.01$. However, it is confirmed later that there exists a Kosterlitz-Thouless (KT) like phase in the ferromagnetic 6-state clock model on the square lattice through studies of the interfacial free energy estimated by Monte Carlo simulations and the authors estimated[10] that lower and upper transition temperatures are 0.75 and 0.90 respectively. Interestingly, in a recent study[11], the machine-learning detection of BKT transition was found in q-state clock model with the neural network based estimates of the two critical points as ${{k_{B}T}\over{J}}=0.410$ and ${{k_{B}T}\over{J}}=0.921$.
The classical 2D clock model is known to have a critical phase with Berezinskii–Kosterlitz–Thouless (BKT) transitions. These transitions have logarithmic corrections which make numerical analysis difficult. Hence, the classical 2D 6-state clock model has been studied [12] by mapping it to the quantum self-dual one dimensional 6-state clock model and confirmed that the self-dual point has a precise numerical agreement with the analytical result. The degenreracy of the excited states at the self dual point has also been arugued from the perspective of the effective field theory.
@@
The thermodynamic behaviours of clock model in layered structured lattices are not yet reported in the well developed literature of this specific model. However, the behaviours of layered Ising and Blume-Capel models are well studied. Let us briefly mention a few selected studies on layered Ising and Blume-Capel models. The main interesting phenomenon in the thermodynamic behaviours of layered Ising and Blume-Capel models is the compensation transition. The compensation temperatures and magnetic susceptibility of a mixed ferro-ferrimagnetic ternary alloy have been studied[18] by Monte Carlo simulations and observed outstanding behaviour of magnetic susceptibility near the compensation and critical temperature. The compensation temperature has been found[19] to depend of the anisotropy in a mixed spin (spin-${1}\over{2}$} and spin-$1$) Heisenberg ferrimagnet on a simple cubic lattice, studied by Oguchi approximation. The anisotopy dependent spin compensation temperature has also been noticed[20] in mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) ferrimagnetic Ising model studied by mean field approximation. In this case, the possibility of multiple compensation temperature was indicated. The impurity- concentration dependent compensation[21] and the anisotropy dependent compensation[22] was also noticed in the trilayered Ising and Blume-Capel models, respectively. The meanfield and effective field theory has been employed [23] to study the interaction dependent compensation in spin-${1}\over{2}$} Ising trilayered ferrimagnetic system. The phase diagram (compensation and no-compensation) was drawn in parameter space. The antiferromagnetic systems showed the compensation behaviours in Ising mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) core/shell nanowire[24] and in mixed spin (spin-${7}\over{2}$} and spin-1{1}) antiferromagnetic ovalene nanowire[25]. The compensation was found[26] Ising mixed spin (spin-${3}\over{2}$} and spin-${1}\over{2}$}) graphene bilayered ferrimagnet. The effects of exchange interactions, crystal field and external magnetic field on the compensation temperature were studied [26] by Monte Carlo simulation. The parameter dependent compensation temperature has been noticed[27] in the Monte Carlo study of mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) Ising model. However, no such compensation was noticed[28] in the study of such kind of mixed spin system in the presence of magnetic field and with ferri and ferromagnetic exchange interactions. The multilayered (composed of two blocks: the upper block A with spin-${5}\over{2}$} and the bottom one B with spin-${3}\over{2}$}) system has been studied[29] by Monte Carlo simulation to investigate the critical and compensation behaviour. The compensation behaviour has been investigated[30] in mixed spin (spin-${3}\over{2}$} and spin-${1}\over{2}$}) Ising ferrimagnetic system with site dilution by nonmagnetic impurities. The compensation temperature has been found to depend strongly on the dilution and interestingly a critical dilution was recorded above which the compensation disappeared.
The evidence of engineering the compensation in ferrimagnetic iron garnet by tuning the ratio of in-plane and out of plane magnetisations can be found in [31]. An interesting spin compensation and first-order reentrant behaviour has been observed[32] in mixed spin (spin-1 and spin-2) Ising ferrimagnetic system studied by Monte Carlo simulation.
@@ 2 Model and numerical details
We consider a classical trilayered (A-B-A type) 6-state clock model defined on a simple cubic lattice, where each of the three layers consists exclusively of either type-A or type-B atoms. These different types of atoms (in each layer) are distinguished by different intralayer ferromagnetic interaction strengths. The magnetic interactions and anisotropic contributions to the energy of the system are captured by the following Hamiltonian:
$\color[rgb]{0,0,1}\mathcal{H}=-J_{ferro}^{AA}\sum_{\langle ij\rangle}\vec{S_{i}}\cdot\vec{S_{j}}-J_{ferro}^{BB}\sum_{\langle ij\rangle}\vec{S_{i}}\cdot\vec{S_{j}}-J_{antiferro}^{AB}\sum_{\langle ij\rangle}\vec{S_{i}}\cdot\vec{S_{j}}-D\sum_{i}[({S_{i}}^{x})^{2}-({S_{i}}^{y})^{2}].$} (1)
@@
The first term in Equation-1 accounts for the intralayer ferromagnetic interactions between nearest neighbour spins $\langle ij\rangle$ within the same atomic plane (for top and bottom layers of A-type atom), with a positive coupling exchange $J_{ferro}^{AA}(>0)$. The ferromagnetic contribution arising from middle layer (B-type of atom) is denoted by the second term, where the nearest neighbour ferromagnetic interaction strength is denoted as $J_{ferro}^{BB}(>0)$. The third term represents interlayer antiferromagnetic coupling between the adjacent layers, involving nearest neighbour sites $\langle ij\rangle$ across the layers, governed by a coupling constant $J_{antiferro}^{AB}(<0)$. The fourth term introduces an anisotropy $D$ (measured in the unit of absolute value of $J_{ferro}^{AB}$) , which destroys the discrete rotational symmetry in the $xy$-plane and favours the spin-alignment in specific direction. Positive values of $D$ prefer specific spin orientations, thereby reinforcing the discrete six-fold symmetry and enhancing the stability of the selected clock states. We impose periodic boundary conditions (PBC) in the x and y directions within each layer and along the stacking direction z, we adopt free boundary conditions to capture the finite thickness of the trilayered structure.
To study the compensation behaviour in the trilayered 6-state clock model, we employ the Monte Carlo simulation technique based on the Metropolis algorithm with random single-spin updates. The system consists of three stacked square lattices, each consisting of $L\times L$ sites, resulting in a total of $3L^{2}$ spins. For the majority of our simulations, we consider a system size of $L=128$, corresponding to $128\times 128$ spins per layer. In order to examine finite-size effects, additional simulations are performed for various lattice sizes (32, 64, 128 and 256).
[…]
To characterize the thermodynamic behaviour of the 6-state clock model, we calculate several key observables that capture the collective magnetic ordering of the system. The sublattice magnetisation provides insight into the local symmetry breaking and is defined for each sublattice $k\in{1,2,3}$ as the average spin components in the x and y directions:
@@
where $\langle E\rangle$ is the average energy density calculated from the Hamiltonian.
@@ 3.1 Isotropic regime:$D=0$
To visualise the spin ordering, we also examine the lattice morphology at representative temperatures. Spin configurations snapshots displayed in Figure-3 reveal a transition from a disordered paramagnetic state at high temperature to a well-aligned domain at low temperature for different layers. At very high temperature $T>>T_{c}$ spin are randomly oriented in all three layers. Six colours denote the six specific clock states. For the isotropic regime $(D=0)$, at high temperature, all six directions are equally preferable. As temperature decreases, the spins at each layer progressively align, leading to the nucleation and growth of the magnetic domain, which marks the onset of long-range order and results a non-zero total magnetisation. Near $T_{comp}$, distinct domain patterns emerge with opposing magnetisation contributions, reflecting the layer-wise compensation behaviour. To provide more systematic and statistically quantiative analysis of the spin orientations near the compensation temperature, the statistical distributions of the angles (orientations) of the spins in different layers are studied and represented by bar graphs in Figure-4, through a comprehensive manner. The normalised spin population (fraction of total spins) is plotted as a function of six discrete clock states ( corresponding spin angle 0, $\pi/3$, $2\pi/3$, $3\pi/3$, $4\pi/3$, $5\pi/3$) for each layer. At high temperature, the spin population is nearly uniform across all angles in all three layers, reflecting an isotropic regime where all spin states are equally probable. As temperature decreases towards the critical value, certain spin angles become more populated, indicating the onset of preferred alignment. Near compensation temperature, layer 1 and 3 exhibit a dominant population at $\theta=2\pi/3(n_{i}=2)$, whereas layer 2 prefers $\theta=5\pi/3(n_{i}=5)$. Importantly, the other spin angles are populated such that net magnetisation of the whole system becomes zero, while each layer individually retains a finite magnetisation. Hence, compensation occurs. Upon further cooling to low temperature, layers 1 and 3 show nearly complete alignment at $\theta=2\pi/3$ while layer 2 remains oriented in $5\pi/3$, resulting nonvanishing total magnetisation. The visual and quantitative results together confirm the layered character of magnetic ordering and the competing interactions that drive the compensation phenomenon. This statistically quantitative representation of the spin orientations is consistent with that represented by qualitative morphological respresentation shown in Figure-3, in the context of compensation.
@@ 3.2 Anisotropic regime: $D>0$
To further elucidate the role of anisotropy in shaping the spin morphology, we also examined the angular distribution of spins for the case of $D=2.0$, where the easy-axis energetically favours spin orientations along $\theta=0$ or $\theta=\pi$. This is shown in Figure-9. At high temperature, the majority of spins in all layers are distributed near $\theta=0(n_{i}=0)$, and $\theta=\pi(n_{i}=3)$ reflecting the dominant easy-axis alignment imposed by the positive anisotropy. Upon cooling, the distribution becomes progressively more ordered, with the relative populations $\theta=0$ and $\theta=\pi$ starting to compete. Near the compensation temperature, a distinct redistribution occurs: layers 1 and 3 exhibit a stronger spin population at $\theta=\pi(n_{i}=3)$, whereas layer 2 retains it alignment at $\theta=0(n_{i}=0)$. This antiparallel arrangement of sublattices leads to a vanishing total magnetisation while preserving finite sublattice magnetisation. At very low temperature, the system reaches a nearly complete alignment where layers 1 and 3 are oriented along $\theta=\pi$ and layer 2 aligned along $\theta=0$. At sufficiently low temperatures, the system recovers a finite total magnetisation as thermal fluctuations subside. Thus, while both the isotropic ($D=0$) and anisotropic ($D=2.0$) systems exhibit a compensation behaviour governed by the redistribution of opposing spin populations, the positive anisotropy merely selects preferred axes ($\theta=0,\pi$) without changing the underlying mechanism.
3.3 Finite size effects:
@@
Finite-size effects play a crucial role in accurately characterising the thermodynamic phase transitions in numerical simulations. To investigate the finite-size behaviour of the clock model in anisotropic regime ($D=2.0$), we performed simulations for different system sizes $L=32,64,128,256$. The magnetisation $M(T)$ and corresponding susceptibility $\chi(T)$ are studied as a function of temperature. The numerical results are shown in Figure-10 for the considered system sizes. Figure-10(a) shows the magnetisation $M(T)$, revealing the expected order-disorder behaviour as the temperature crosses the transition ($T_{c}$) and the low temperature compensation. The susceptibility curves for different system sizes (shown in Figure-10(b))manifest two salient features: a prominent peak at the thermodynamic critical temperature and a secondary smeared peak at the compensation point. The height of the peak of the susceptibility at $T_{c}$ exhibits a systematic increase with increasing system sizes, confirming the growth of critical correlation for the second-order phase transition. The susceptibility is believed to be divergent in the thermodynamic limit ($L\rightarrow\infty$). In contrast, the secondary smeared peak associated with the compensation point does not show such behaviour; its height remains nearly constant across the studied system sizes. This suggests that such compensation is not a critical transition, but rather a finite, non-divergent reorganisation of the magnetisation in the magnetic sublattices. The finite-size analysis confirms that the clock model undergoes a thermodynamic phase transition at $T_{c}$ , where critical fluctuations become system-size dependent, while the compensation point reflects a non-critical feature, inherent to the system’s response.